how would you divide this on paper? 12 divide by .15

Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12844)

12/.15 REDUCE BY DIVIDING BY 3.
4/.05 MULTIPLY BY 100
400/5 DIVIDE.
80 ANS.

Answer by richwmiller(17219)

.15=15/100
multiplying by the reciprocal is the same as division
12*100/15=4/5*100=80